How many 6 digit numbers can be made with any combination and permutation of 3 and 7? (e.g. 377777, 737777, 777333, 733733, 333337): How many 6 digit numbers can be made with any combination and permutation of 3 and 7? (e.g. 377777, 737777, 777333, 733733, 333337)
Answer by jim_thompson5910(9929)

Let's try to find the number of combinations that 3 and 7 can make up for a two digit number.

Let's try a smaller example. So let's use the numbers 3 and 7 to make a two digit number, three digit number, and a four digit number to see what is going on


So let's list ALL of the possible two digit numbers:

33
37
73
77

So for a two digit number, there are 4 possible combinations


-----------------------------

So let's list ALL of the possible three digit numbers:

333
337
373
377
733
737
773
777

So for a three digit number, there are 8 different combinations.
-----------------------------

So let's list ALL of the possible four digit numbers:

3333
3337
3373
3377
3733
3737
3773
3777
7333
7337
7373
7377
7733
7737
7773
7777

So for a four digit number, there are 16 different combinations.


Notice how the number of combinations keeps doubling (since 2 doubles to 4, 4 doubles to 8, 8 doubles to 16, etc...).

So if we write out a 5 digit number only using 3 and 7, then we will list out 32 different combinations (which is double of 16). And finally for a six digit number, there will be 64 different combinations (which is double of 32).


Note: the quicker way (ideally once you understand what's going on) is to use this formula y=2^x where x is the number of digits and y is the number of unique number combinations.

So in this case, x=6 and y=2^6=64. So once again, we get 64 different combinations.


--------------------------------------------------

Answer:

So there are 64 unique 6 digit numbers that can be generated from just the two values 3 and 7.